# The Connection between Dusty Star-Forming Galaxies and the First Massive Quenched Galaxies

### v1

@@
11institutetext: Departamento de Astronomia, Instituto de Astronomia, Geofísica e Ciências Atmosféricas, Universidade de São Paulo, Rua do Matão 1226, Cidade Universitária, 05508-900, São Paulo, SP, Brazil 22institutetext: Cosmic Dawn Center (DAWN), Copenhagen, Denmark 33institutetext: DTU Space, Technical University of Denmark, Elektrovej 327, DK2800 Kgs. Lyngby, Denmark 44institutetext: Jodrell Bank Centre for Astrophysics, University of Manchester, Oxford Road, Manchester M13 9PL, UK 55institutetext: Centre for Astrophysics Research, University of Hertfordshire, Hatfield, AL10 9AB, UK 66institutetext: Eureka Scientific, Inc., 2452 Delmer Street, Suite 100, Oakland, CA 94602, USA 77institutetext: Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, The University of Tokyo, Kashiwa, Chiba 277-8583, Japan 88institutetext: Leiden Observatory, Leiden University, PO Box 9513, NL-2300 RA Leiden, the Netherlands
@@ 1 Introduction
The mechanisms driving the early quenching of high-redshift MQs remain unclear. X-ray and radio emission have been detected in a subset of MQs (Marsan et al., 2017; Ito et al., 2022; Kubo et al., 2022). Notably, MQs tend to host stronger AGN activity — as indicated by both X-ray and radio luminosities — than star-forming galaxies of similar stellar mass, whose relative excess increases with redshift (Ito et al., 2022). This supports the view that AGN feedback may be the dominant quenching mechanism for MQs at $z\gtrsim 3$. On the other hand, other studies (e.g., Ellison et al., 2022, 2024; Calderón-Castillo & Smith, 2024; Gordon et al., 2025) suggest that mergers primarily deplete galaxies’ cold-gas reservoirs through intense starburst episodes, rather than through AGN-driven feedback. However, the relatively short detectability timescales of AGN ($\Delta t\sim 10^{5}\,$yr; Schawinski et al., 2015) may result in the underestimation of the true prevalence of recent AGN activity in quenching galaxies. MQs have also been identified in protocluster environments at $z\sim 3$; the results of Kubo et al. (2021), McConachie et al. (2022), and Ito et al. (2023) support a scenario in which these galaxies quenched due to accelerated internal evolution — such as elevated star formation rates combined with AGN and/or SN feedback — rather than explicit environmental effects such as ram pressure stripping. In contrast, Shi et al. (2021) reported evidence for environmental quenching in the D4UD01 protocluster at $z=3.24$, suggesting that multiple quenching pathways may coexist.
@@
While MQs are rare, with number densities of $\sim 10^{-5}\ \mathrm{Mpc}^{-3}$ at $z\sim 3.5$ (Schreiber et al., 2018; Straatman et al., 2014; Merlin et al., 2019) (though estimates vary significantly between studies; see the compilation presented by Valentino et al. 2023), their number densities at high redshift are in tension with predictions from most current galaxy formation models. Accurately reproducing the observed abundance of MQs at $z\gtrsim 3$ at the same time as matching low-redshift constraints has proven particularly challenging; this tension has only been exacerbated by recent JWST results, which suggest an even more abundant MQ population than was inferred pre-JWST. A recent systematic comparison by Lagos et al. (2025) assessed the success of various semi-analytic models (SAMs) — including GAEA (De Lucia et al., 2024), GALFORM (Lacey et al., 2016), and SHARK v2.0 (Lagos et al., 2024) — as well as hydrodynamical simulations such as SIMBA, IllustrisTNG, and EAGLE, in reproducing the number densities of MQs at $z\gtrsim 3$. Lagos et al. (2025) showed that most of these models under-predict the observationally-inferred MQ number densities at $z\sim 4$, by between $0.3\,\rm{dex}$ and more than an order of magnitude. Similar discrepancies have been reported by other groups — for example, Szpila et al. (2025) using the SIMBA-C simulation (Hough et al., 2023, 2024), and Vani et al. (2025) using an updated version of the L-Galaxies SAM (Ayromlou et al., 2021).
@@
In this work, we adopt the Planck Collaboration et al. (2014) cosmology: $\sigma_{8}=0.829$, $H_{0}=67.3\,\rm{km\,s}^{-1}\,\rm{Mpc}^{-1}$, $\Omega_{\Lambda}=0.685$, $\Omega_{m}=0.315$, $\Omega_{b}=0.0487$, $f_{b}=0.155$, and $n=0.96$, consistent with the cosmologically rescaled version of the Millennium simulation (Angulo & Hilbert, 2015). Throughout this study, we assume a Chabrier (2003) IMF.
@@ 2 Galaxy Formation Model
We use the Henriques et al. (2020) version of the L-Galaxies semi-analytical model (SAM) of galaxy formation, specifically the re-calibrated version from Araya-Araya et al. (2025). We run the model on the Millennium simulation (Springel, 2005), scaled to the Planck Collaboration et al. (2014) cosmology using the method from Angulo & Hilbert (2015). This scaled version of the Millennium simulation has a total box volume of $(713.6\,\rm{cMpc})^{3}$ and a dark matter particle mass resolution of $m_{p}=1.43\times 10^{9}\,\rm{M_{\odot}}$ ($h=0.673$, in Planck Collaboration et al. (2014) cosmology). In this section, we describe the key features of the galaxy formation model.
The merger trees of dark matter halos, constructed with the SUBFIND algorithm (Springel et al., 2001), form the backbone of L-Galaxies. These virialised halos are populated with baryonic matter, beginning with primordial diffuse gas. Initially, the gas infalls into the halos, forming a hot atmosphere. Upon cooling, this gas settles into a disk, where it becomes available for star formation, regulated by the surface density of molecular hydrogen ($\rm{H}_{2}$). Over time, stars die and release energy, enriching the interstellar medium (ISM) and circumgalactic medium (CGM) with metals. The energy from supernovae (SNe) can (1) heat the cold gas in the disk, (2) reheat gas in the hot atmosphere—suppressing further cooling, and (3) drive outflows that move gas to an external reservoir, from which it may later re-accrete. The Henriques et al. (2020) version of the model also includes metal enrichment from AGB stars, as well as from SN-Ia and SN-II events. All these processes are tracked in concentric rings within galaxies.
@@
Table 1: A summary of the L-Galaxies model used in this paper. These best-fit parameters were obtained by Araya-Araya et al. (2025) for the ‘no HIMF’ model, which obtains a good match to the $S_{870}$ number counts while remaining consistent with observationally-derived lower limits on the number density of high-redshift massive quiescent galaxies. The model was calibrated using the MCMC mode of L-Galaxies. The observational constraints were the stellar mass function and fraction of quiescent ($\mathrm{sSFR}\leq 10^{-11}\,\mathrm{yr}^{-1}$) galaxies as a function of stellar mass at $z=0.4$ and $z=2.8$ (Leja et al., 2020, 2022), and the number density of bright ($S_{870}\geq 5.2\,\mathrm{mJy}$) sub-millimetre galaxies at $z=2.8$ (Dudzevičiūtė et al., 2020). Full details of these parameters and their corresponding equations are given in the Supplementary Material of Henriques et al. (2020).
@@
Figure 1: The star formation rate (SFR)–stellar mass ($M_{\star}$) relation at $z=2$, 3, and 4 is shown in the first, second, and third panels, respectively, using the re-calibrated version of L-Galaxies from Araya-Araya et al. (2025). The coloured dots represent bright DSFGs with $S_{870}\geq 1\,\rm{mJy}$, colour-coded by their $S_{870}$ flux density (see colour-bar), derived using the scaling relations presented by Cochrane et al. (2023b). The red dashed lines indicate the thresholds in specific star formation rate (sSFR) and stellar mass used to separate quenched and massive galaxies, respectively, following the definition of Carnall et al. (2020). The pink line and colored area correspond to the ‘main sequence’ of galaxies in our model and its $1\sigma$ dispersion. Overall, bright DSFGs are located in the massive, star-forming region of the diagram and lie on, or just above, the ‘main sequence’. As expected, the massive, quenched region is less populated at higher redshifts.
@@ 3 Results
Figure 2: The evolution of the $S_{870}$ flux density of massive quiescent galaxies (MQs) selected at $z=2$, 3, and 4 is shown in the first, second, and third panels, respectively. At each selected redshift, MQs are divided into four stellar mass bins of width $0.4\,\rm{dex}$, starting from a lower limit of $\log(M_{\star}/\rm{M_{\odot}})=10.6$. The colored regions and lines represent the $16^{\rm{th}}$, $50^{\rm{th}}$, and $84^{\rm{th}}$ percentiles of MQs within each stellar mass bin. Overall, the $S_{870}$ evolution of MQs exhibits a gradual increase with decreasing redshift, reaching a peak before undergoing a rapid decline at later times. Notably, the most massive MQs consistently display the highest $S_{870}$ peak values.
@@ 3.1 High-redshift Massive Quiescent Galaxies were sub-millimetre bright at earlier cosmic times
As described above, we select massive MQs at $z=2$, 33, and $4$. In order to analyse the evolution of these selected galaxies across cosmic time, we track each galaxy along its so-called Main Progenitor Branch (MPB) by matching galaxies at different redshifts with the same MainLeafID (i.e. that have the same original progenitor galaxy). For simplicity, we do not follow galaxies once they are accreted by a more massive system. However, the fraction of MQs that are eventually incorporated into a more massive galaxy is modest ($\lesssim 24\%$), regardless of the selected redshift.
First, we explore the past evolution of sub-millimetre flux density for our MQ samples selected at different redshifts. In Figure 2, we show the median evolution of $S_{870}$ flux density for MQs selected at $z=2$, 33, and $4$ in four stellar mass (at the selected redshift) bins. The most notable trend in Figure 2 is that MQs tend to reach a peak in sub-millimetre flux density before the redshift at which they are selected. The typical $S_{870}$ evolution exhibits a gradual increase with cosmic time (i.e. with decreasing redshift), between $z\sim 9$ and its peak, followed by a rapid decline at lower redshifts. This can be seen for MQs at the three analyzed redshifts. Another noteworthy trend is that, on average, MQs with higher stellar masses tend to reach higher peak sub-millimetre flux densities.
@@
Another tentative correlation apparent in Figure 2 is that more massive MQs tend to have reached higher peak sub-millimetre fluxes. To explore this further, we present in Figure 4 the maximum $870\,\mu\rm{m}$ flux density reached by MQs ($S_{870,{\rm max}}$) as a function of their stellar mass at the selected redshift. Our results reveal a strong correlation between these two properties across all three redshifts analysed. The scatter in this relation is typically less than $\sim 0.5\,\mathrm{dex}$ and tends to decrease with increasing stellar mass. Interestingly, at fixed stellar mass, the deviation from the median relation is correlated with the redshift at which the sub-millimetre peak occurred: MQs with a given stellar mass, $M_{\star}$, reached a brighter $S_{870,{\rm max}}$ if the sub-millimetre emission peaked at an earlier epoch.
@@
Table 2: The best-fit parameters for Equation 2, which relates the maximum sub-millimetre flux-density reached by massive quiescent galaxies selected at $z({\rm MQ})$, across their history ($S_{870,{\rm max}}$) to the stellar mass at $z({\rm MQ})$.
@@ 3.2 Not All Dusty Star-Forming Galaxies Are Progenitors of High-redshift MQs
We have already shown that, in our model, massive quiescent galaxies at $z=2$, 3, and 4 were sub-millimetre bright in the past. For instance, we find that $\sim 86\%$, $96\%$, and $91\%$ of MQs at $z=2$, 3, and 4, exceeded $1\,{\rm mJy}$ at some point in the past. Only $\sim 7\%$, $9\%$, and $5\%$ reached $S_{870,{\rm max}}\geq 3\,{\rm mJy}$, while $\sim 2\%$, $1\%$, and $0\%$ surpassed $5\,{\rm mJy}$. However, it is not yet clear whether a significant fraction of all DSFGs – defined here as galaxies with $S_{870}\geq 1\,{\rm mJy}$ – at those redshifts are indeed progenitors of MQs at $2\leq z\leq 4$.
@@
To investigate the primary physical processes responsible for quenching star formation in DSFG and MQ progenitors, we explore the evolution of the median energy injection rates due to supernova (SN) and AGN feedback for the four galaxy populations at the three selected redshifts, as shown in Figure 8. The most prominent trend in Figure 8 is that the SN feedback energy injection rate gradually increases from early times ($z\sim 9$), peaking at a redshift that coincides with the onset of AGN feedback — which also matches the average redshift of the first relevant merger (see Figure 7). After this peak, the SN energy input rapidly declines, a trend that can be attributed to the drop in star formation following the intense merger-induced starburst, particularly in MQ progenitors.

### v2

@@
11institutetext: Cosmic Dawn Center (DAWN), Copenhagen, Denmark 22institutetext: DTU Space, Technical University of Denmark, Elektrovej 327, DK2800 Kgs. Lyngby, Denmark 33institutetext: Departamento de Astronomia, Instituto de Astronomia, Geofísica e Ciências Atmosféricas, Universidade de São Paulo, Rua do Matão 1226, Cidade Universitária, 05508-900, São Paulo, SP, Brazil 44institutetext: Jodrell Bank Centre for Astrophysics, University of Manchester, Oxford Road, Manchester M13 9PL, UK 55institutetext: Centre for Astrophysics Research, University of Hertfordshire, Hatfield, AL10 9AB, UK 66institutetext: J.P. Morgan Securities LLC, 390 Madison Avenue, New York, NY 10017, USA 77institutetext: Leiden Observatory, Leiden University, PO Box 9513, NL-2300 RA Leiden, the Netherlands
@@ 1 Introduction
The mechanisms driving the early quenching of high-redshift MQs remain unclear. X-ray and radio emission have been detected in a subset of MQs (Marsan et al., 2017; Ito et al., 2022; Kubo et al., 2022). Notably, MQs tend to host stronger AGN activity — as indicated by both X-ray and radio luminosities — than star-forming galaxies of similar stellar mass, whose relative excess increases with redshift (Ito et al., 2022). This supports the view that AGN feedback may be the dominant quenching mechanism for MQs at $z\gtrsim 3$. On the other hand, other studies (e.g., Ellison et al., 2022, 2024; Calderón-Castillo and Smith, 2024; Gordon et al., 2026) suggest that mergers primarily deplete galaxies’ cold-gas reservoirs through intense starburst episodes, rather than through AGN-driven feedback. However, the relatively short detectability timescales of AGN ($\Delta t\sim 10^{5}\,$yr; Schawinski et al., 2015) may result in the underestimation of the true prevalence of recent AGN activity in quenching galaxies. MQs have also been identified in protocluster environments at $z\sim 3$; the results of Kubo et al. (2021), McConachie et al. (2022), and Ito et al. (2023) support a scenario in which these galaxies quenched due to accelerated internal evolution — such as elevated star formation rates combined with AGN and/or SN feedback — rather than explicit environmental effects such as ram pressure stripping. In contrast, Shi et al. (2021) reported evidence for environmental quenching in the D4UD01 protocluster at $z=3.24$, suggesting that multiple quenching pathways may coexist.
@@
While MQs are rare, with number densities of $\sim 10^{-5}\ \mathrm{Mpc}^{-3}$ at $z\sim 3.5$ (Schreiber et al., 2018; Straatman et al., 2014; Merlin et al., 2019) (though estimates vary significantly between studies; see the compilation presented by Valentino et al. 2023), their number densities at high redshift are in tension with predictions from most current galaxy formation models. Accurately reproducing the observed abundance of MQs at $z\gtrsim 3$ at the same time as matching low-redshift constraints has proven particularly challenging; this tension has only been exacerbated by recent JWST results, which suggest an even more abundant MQ population than was inferred pre-JWST. A recent systematic comparison by Lagos et al. (2025) assessed the success of various semi-analytic models (SAMs) — including GAEA (De Lucia et al., 2024), GALFORM (Lacey et al., 2016), and SHARK v2.0 (Lagos et al., 2024) — as well as hydrodynamical simulations such as SIMBA, IllustrisTNG, and EAGLE, in reproducing the number densities of MQs at $z\gtrsim 3$. Lagos et al. (2025) showed that most of these models under-predict the observationally-inferred MQ number densities at $z\sim 4$, by between $0.3\,\rm{dex}$ and more than an order of magnitude. Similar discrepancies have been reported by other groups — for example, Szpila et al. (2025) using the SIMBA-C simulation (Hough et al., 2023, 2024), and Vani et al. (2025) using an updated version of the L-Galaxies SAM (Ayromlou et al., 2021). On the other hand, the observed number densities of MQs at high redshift could be significantly contaminated, for example by dusty galaxies or ‘Little Red Dots’ (Yang et al., 2025), a population of compact, presumably AGN-dominated sources (Labbé et al., 2023; Greene et al., 2024; Matthee et al., 2024; Li et al., 2025), recently uncovered by JWST observations.
@@
In this work, we adopt the Planck Collaboration et al. (2014) cosmology: $\sigma_{8}=0.829$, $H_{0}=67.3\,\rm{km\,s}^{-1}\,\rm{Mpc}^{-1}$, $\Omega_{\Lambda}=0.685$, $\Omega_{m}=0.315$, $\Omega_{b}=0.0487$, $f_{b}=0.155$, and $n=0.96$, consistent with the cosmologically rescaled version of the Millennium simulation (Angulo and Hilbert, 2015). Throughout this study, we assume a Chabrier (2003) IMF.
@@ 2 Galaxy formation model
We use the Henriques et al. (2020) version of the L-Galaxies semi-analytical model (SAM) of galaxy formation, specifically the re-calibrated version from Araya-Araya et al. (2025). We run the model on the Millennium simulation (Springel, 2005), scaled to the Planck Collaboration et al. (2014) cosmology using the method from Angulo and Hilbert (2015). This scaled version of the Millennium simulation has a total box volume of $(713.6\,\rm{cMpc})^{3}$ and a dark matter particle mass resolution of $m_{p}=1.43\times 10^{9}\,\rm{M_{\odot}}$ ($h=0.673$, in Planck Collaboration et al. (2014) cosmology). In this section, we describe the key features of the galaxy formation model.
The merger trees of dark matter halos, constructed with the SUBFIND algorithm (Springel et al., 2001), form the backbone of L-Galaxies. These virialised halos are populated with baryonic matter, beginning with primordial diffuse gas. Initially, the gas infalls into the halos, forming a hot atmosphere. Upon cooling, this gas settles into a disk, where it becomes available for star formation, regulated by the surface density of molecular hydrogen ($\rm{H}_{2}$). Over time, stars die and release energy, enriching the interstellar medium (ISM) and circumgalactic medium (CGM) with metals. The energy from supernovae (SNe) can (1) heat the cold gas in the disk, moving it to hot atmosphere, available for further cooling, and (2) drive outflows that move gas to an external reservoir, from which it may later re-accrete. The Henriques et al. (2020) version of the model also includes metal enrichment from AGB stars, as well as from SN-Ia and SN-II events. All these processes are tracked in concentric rings within galaxies.
@@
Table 1: Summary of the L-Galaxies model used in this paper. [^2] These best-fit parameters were obtained by Araya-Araya et al. (2025) for the ‘no HIMF’ model, which obtains a good match to the $S_{870}$ number counts while remaining consistent with observationally-derived lower limits on the number density of high-redshift massive quiescent galaxies. The model was calibrated using the MCMC mode of L-Galaxies. The observational constraints were the stellar mass function and fraction of quiescent ($\mathrm{sSFR}\leq 10^{-11}\,\mathrm{yr}^{-1}$) galaxies as a function of stellar mass at $z=0.4$ and $z=2.8$ (Leja et al., 2020, 2022), and the number density of bright ($S_{870}\geq 5.2\,\mathrm{mJy}$) sub-millimetre galaxies at $z=2.8$ (Dudzevičiūtė et al., 2020). Full details of these parameters and their corresponding equations are given in the Supplementary Material of Henriques et al. (2020).
@@
Figure 1: The star formation rate (SFR)–stellar mass ($M_{\star}$) relation at $z=2$, 3, and 4 is shown in the first, second, and third panels, respectively, using the re-calibrated version of L-Galaxies from Araya-Araya et al. (2025). The coloured dots represent bright DSFGs with $S_{870}\geq 1\,\rm{mJy}$, colour-coded by their $S_{870}$ flux density (see colour-bar), derived using the scaling relations presented by Cochrane et al. (2023b). The red dashed lines indicate the thresholds in specific star formation rate (sSFR) and stellar mass used to separate quenched and massive galaxies, respectively, following the definition of Carnall et al. (2020). The pink line and coloured area correspond to the ‘main sequence’ of galaxies in our model and its $1\sigma$ dispersion. Overall, bright DSFGs are located in the massive, star-forming region of the diagram and lie on, or just above, the ‘main sequence’. As expected, the massive, quenched region is less populated at higher redshifts.
@@ 3 Results
Figure 2: The evolution of the $S_{870}$ flux density of massive quiescent galaxies (MQs) selected at $z=2$, 3, and 4 is shown in the first, second, and third panels, respectively. At each selected redshift, MQs are divided into four stellar mass bins of width $0.4\,\rm{dex}$, starting from a lower limit of $\log(M_{\star}/\rm{M_{\odot}})=10.6$. The coloured regions and lines represent the $16^{\rm{th}}$, $50^{\rm{th}}$, and $84^{\rm{th}}$ percentiles of MQs within each stellar mass bin. Overall, the $S_{870}$ evolution of MQs exhibits a gradual increase with decreasing redshift, reaching a peak before undergoing a rapid decline at later times. Notably, the most massive MQs consistently display the highest $S_{870}$ peak values.
@@ 3.1 High-redshift massive quiescent galaxies were sub-millimetre bright at earlier cosmic times
As described above, we select massive MQs at $z=2$, 33, and $4$. In order to analyse the evolution of these selected galaxies across cosmic time, we track each galaxy along its so-called Main Progenitor Branch (MPB) by matching galaxies at different redshifts with the same MainLeafID (i.e. that have the same original progenitor galaxy). First, we explore the past evolution of sub-millimetre flux density for our MQ samples selected at different redshifts. In Figure 2, we show the median evolution of $S_{870}$ flux density for MQs selected at $z=2$, 33, and $4$ in four stellar mass (at the selected redshift) bins. The most notable trend in Figure 2 is that MQs tend to reach a peak in sub-millimetre flux density before the redshift at which they are selected. The typical $S_{870}$ evolution exhibits a gradual increase with cosmic time (i.e. with decreasing redshift), between $z\sim 9$ and its peak, followed by a rapid decline at lower redshifts. This can be seen for MQs at the three analysed redshifts. Another noteworthy trend is that, on average, MQs with higher stellar masses tend to reach higher peak sub-millimetre flux densities.
@@
Another tentative correlation apparent in Figure 2 is that more massive MQs tend to have reached higher peak sub-millimetre fluxes. To explore this further, we present in Figure 4 the maximum $870\,\mu\rm{m}$ flux density reached by MQs ($S_{870,{\rm max}}$) as a function of their stellar mass at the selected redshift. Our results reveal a strong correlation between these two properties across all three redshifts analysed. The scatter in this relation is typically less than $\sim 0.5\,\mathrm{dex}$ and tends to decrease with increasing stellar mass. Interestingly, at fixed stellar mass, the deviation from the median relation is correlated with the redshift at which the sub-millimetre peak occurred: MQs with a given stellar mass, $M_{\star}$, reached a brighter $S_{870,{\rm max}}$ if the sub-millimetre emission peaked at an earlier epoch. This correlation arises because, in our model, galaxies that reached their sub-millimetre peak earlier have, on average, higher star-formation rates than those at lower redshifts.
@@
Table 2: The best-fit parameters for Equation 2333Equation 2 relates the maximum sub-millimetre flux-density reached by massive quiescent galaxies selected at $z({\rm MQ})$, across their history ($S_{870,{\rm max}}$) to the stellar mass at $z({\rm MQ})$.
@@ 3.2 Not all dusty star-forming galaxies are progenitors of high-redshift MQs
We have already shown that, in our model, massive quiescent galaxies at $z=2$, 3, and 4 were sub-millimetre bright in the past. For instance, we find that $\sim 86\%$, $96\%$, and $91\%$ of MQs at $z=2$, 3, and 4, exceeded $1\,{\rm mJy}$ at some point in the past. Only $\sim 7\%$, $9\%$, and $5\%$ reached $S_{870,{\rm max}}\geq 3\,{\rm mJy}$, while $\sim 2\%$, $1\%$, and $0\%$ surpassed $5\,{\rm mJy}$. However, it is not yet clear whether a significant fraction of all DSFGs – defined here as galaxies with $S_{870}\geq 1\,{\rm mJy}$ – at those redshifts are indeed progenitors of MQs at $2\leq z\leq 4$. For simplicity, we do not follow galaxies once they are accreted by a more massive system. However, the fraction of MQs that are eventually incorporated into a more massive galaxy is modest ($\lesssim 24\%$), regardless of the selected redshift.
@@
To investigate the primary physical processes responsible for quenching star formation in DSFG and MQ progenitors, we explore the evolution of the median energy injection rates due to supernova (SN) and AGN feedback444see Appendix A for a detailed description of how these quantities are calculated. for the four galaxy populations at the three selected redshifts, as shown in Figure 8. The most prominent trend in Figure 8 is that the SN feedback energy injection rate gradually increases from early times ($z\sim 9$), peaking at a redshift that coincides with the onset of AGN feedback — which also matches the average redshift of the first relevant merger (see Figure 7). After this peak, the SN energy input rapidly declines, a trend that can be attributed to the drop in star formation following the intense merger-induced starburst, particularly in MQ progenitors.